You want to create a portfolio equally as risky as the market, and you have $1,800,000 to invest. Given this information, fill in the rest of the following table: (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations.)

asset investment beta
stock A $396,000 1.20
stock B $504,000 1.40
stock C 1.60
risk free asset
Answer #1

Let investment in C=$x

Hence investment in risk free asset=1,800,000-(396,000+504,000+x)=$(900,000-x)

Portfolio beta=Respective beta*Respective weight

1=(396,000/1,800,000*1.2)+(504,000/1,800,000*1.4)+(x/1,800,000*1.6)+(900,000-x)/1,800,000*0[Beta of market=1;Beta of risk-free assets=0]

1=0.656+(x/1,800,000*1.6)

x=(1-0.656)*1,800,000/1.6

=$387,000=investment in C

Hence investment in risk free asset=$(900,000-x)

=$513,000

Beta of risk free assets=0